OpenStudy (anonymous):

Select the best possible first step to solving the system by first eliminating the x variable. 3x + 5y = 1 2x + 4y = -4 A. Multiply the first equation by 2 and multiply the second equation by 3. B.Multiply the first equation by -2 and multiply the second equation by 3. C.Multiply the first equation by -2 and multiply the second equation by -3. D.None of the above

OpenStudy (anonymous):

To eliminate x you first gotta multiply the equation 2 by the equation 1.So it will means that 3x+5y=1(the first equation) will be multiplied by 2.And for the second equation you gotta multiply 3 times 2x+4y=-4(second equation.)

OpenStudy (anonymous):

GRACIASSSSSSS !!!!! @joshhfers

OpenStudy (anonymous):

By any chance you have any idea for this one How can x2 = x2 + 2x + 9 be set up as a system of equations? :/ ?